OpenStudy (anonymous):

Can someone work with me to check some of my homework answers? Suppose that the functions g and h are defined for all real numbers x as follows. g(x) = x – 2 h(x) = 4x + 2 Write the expressions for (g * h) and (g – h ) (x) and evaluate (g + h) (1). My answers: (g * h) (x) = 64 (g – h) (x) = 4x^2 – 6x – 4 (g + h) (s) = 4x^2 + 6x + 4

OpenStudy (anonymous):

I believe you are correct. But hold on.

OpenStudy (anonymous):

@blurbendy

OpenStudy (anonymous):

Ok.

OpenStudy (blurbendy):

i dont see how you got 64, for g(x) * h(x)

OpenStudy (anonymous):

Ok, what did you get?

OpenStudy (blurbendy):

(x-2)(4x +2) then foil

OpenStudy (blurbendy):

= 4x^2 -6x -4

OpenStudy (blurbendy):

g - x = (x -2) - (4x + 2) = -3x -4

OpenStudy (blurbendy):

g - h, not g - x, oops

OpenStudy (anonymous):

When I multiplied (g * h) (x) = 64 I got this answer. Are you saying all of the answers are incorrect?

OpenStudy (anonymous):

Can you tell me what the correct answers are?

OpenStudy (loser66):

read blurbendy comments, he gave you all of the answers.